# Learning from Imperfect Human Feedback: a Tale from Corruption-Robust $\text{Dueling}^{*}$

### v1

@@ C.4 Proof for Proposition 2: Efficiency-Robustness Tradeoff in NC-SMD
Choosing $\eta_{\rho}=\frac{\sqrt{\log T}}{2d}T^{-\alpha},\alpha\in[0.5,1]$, we have
@@
$\text{Reg}_{T}$ $\leq 2dC_{0}\sqrt{\log T}T^{\alpha}+4d\sqrt{\log T}T^{1-\alpha}+4RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+4R\sqrt{d}L_{\sigma}(\log T)^{0.25}T^{\frac{1}{2}(1-\alpha)}C+2LL_{0}R$
@@
$\leq O\left(d\sqrt{\log T}T^{\alpha}+\sqrt{d}(\log T)^{\frac{1}{4}}T^{\frac{1}{2}(1-\alpha)}C+dC\right),$
which completes the proof.
D Additional Experiments
@@
In this section, we list all the experiment details. At the beginning, we introduce the regret order fitting method, different types of corruption that we simulate in the experiments, and baseline algorithms for comparison.
@@
Corruption. In the experiments, we consider two types of corruption: corruption induced by $\rho$-imperfect user and arbitrary corruption. For corruption induced by $\rho$-imperfect user, we simulate $c_{t}(a_{t},a^{\prime}_{t})$ according to Definition A.1. We set $\lambda=2$ and $C_{0}=0.1(\max_{a\in\mathcal{A}}\mu(a)-\min_{a\in\mathcal{A}}\mu(a))$. Moreover, we add $c_{t}(a_{t},a^{\prime}_{t})$ to the utility difference of $a_{t}$ and $a^{\prime}_{t}$ adversarially. In particular, if $\mu(a_{t})>\mu(a^{\prime}_{t})$, then ^$\hat{\mathbb{P}}(a_{t}\succ a^{\prime}_{t})=\sigma(\mu(a_{t})-\mu(a^{\prime}_{t})-c_{t}(a_{t},a^{\prime}_{t}))$, then vice versa. For arbitrary corruption, we force the user to submit her least preferred item each round to the algorithm over the first $C$ rounds.
Baseline Algorithms. We consider three baseline algorithms for comparison, Doubler (Ailon et al., 2014), Sparring (Ailon et al., 2014), and Versatile-DB (Saha and Gaillard, 2022).
Doubler is the first approach that transforms a dueling bandit problem into a standard multi-armed bandit (MAB) problem. It operates in epochs of exponentially increasing length: in each epoch, the left arm is sampled from a fixed distribution, and the right arm is selected using a MAB algorithm to minimize regret against the left arm. The feedback received by the MAB algorithm is the number of wins the right arm achieves compared to the left arm. Under linear link assumption, Doubler has been proven to experience regret as the same order as underlying MAB algorithm. For continuous action space and general concave utility, we choose Bandit Gradient Descent (BGD, (Flaxman et al., 2004)), with regret $O(T^{3/4})$, as the underlying MAB algorithm.
@@
Versatile-DB applies novel reduction approach which converts dueling bandit to MAB. Essentially, it designs simplex over $K$ arms using the idea of follow-the-regularized-leader and updates the simplex using information of whether the arm wins. It has been proved that Versatile-DB is robust to corruption defined as the number of flips of duels and Versatile-DB enjoys regret linear in $C$ (Saha and Gaillard, 2022).
D.1 Experiments on Synthetic Data
D.1.1 DBGD Lower Bound Simulation
Experiment Setup. To validate Corollary 1, we construct the following problem instance. Consider the linear utility function $\mu(a):=\theta^{\top}a$. Choose $d=2,\theta=[\frac{1}{2},\frac{1}{2}],T=10^{5}$. Let action set $\mathcal{A}:=\{(a_{1},a_{2}):a_{1}\geq 0,a_{2}\geq 0,\frac{1}{2}a_{1}+a_{2}-\frac{1}{4}\leq 0\}$. We repeat the experiments for $50$ times, each time with a different seed.
Results and Discussion. Each grey-colored dotted line in Figure 2 indicate the cumulative regret for each simulation. The blue line is the average cumulative regret over 50 repeats. The error bar indicates $\pm$ one standard deviation of the regret. The fitted order over the last $1\%$ of the average cumulative regret is 0.76, which is close $0.75$ and supports Corollary 1.
D.1.2 Regret Upper Bound in Theorem 1
Experiment Setup. To validate regret upper bound in Theorem 1, we consider a strongly concave utility $\mu_{\theta}(a):=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, a logistic link function $\sigma(x)=\frac{1}{1+\exp(-x)}$. We choose $d=5$, and $T=10^{5}$. Our action space $\mathcal{A}$ is a $d$-dimensional ball with radius $R=10$. The preference parameter $\theta$ is randomly sampled from the surface of $\mathcal{A}$. In our problem setting, the optimal action $a^{*}$ is $\theta$ and $\mu_{\theta}(a^{*})=50$. We simulate corruption induced by $\rho$-imperfect user and conduct experiments for $\rho\in[0.5,1.0]$. For each value of $\rho$, we repeat the experiments for $5$ times, each under a different seed. We set the learning rate $\eta_{\rho}$ according to Proposition C.1.
Figure D.1: Performance under Known C
Results and Discussion. Figure D.1 shows that when $\rho$ is known, NC-SMD can tolerate linear $\rho$-imperfect user corruption, which confirms theoretical analysis in Proposition C.1 and implies that lower bound in Theorem 1 is tight.
D.1.3 Robustness under Corruption induced by Imperfect User
In Figure D.2, we simulated the performance of DBGD, NC-SMD, Sparring, Doubler in the same experiment setting as in Section 6.1 expect using corruption induced by $\rho$-Imperfect User. We notice that the experiment results also align with theoretical prediction (Theorem 2 and Proposition 2), which predicts when $\alpha=0.25$ for DBGD and $\alpha=0.5$ for NC-SMD, they can tolerate $O(T^{0.75})$ agnostic corruption.
Figure D.2: Robustness under Corruption induced by $\rho$-Imperfect User
D.1.4 Efficiency-Robustness Tradeoff in Theorem 2
@@
To validate efficiency-robustness trade-off in Theorem 2 and Proposition 2, we adopt the same experiment setup as in Section 6.1 except different $\alpha$ values. In the first row of Figure D.3, we consider $\alpha=0.05$ for DBGD and $\alpha=0.9$ for NC-SMD. We consider $\rho\in[0.5,0.95]$. According to the theoretical prediction, both DBGD and NC-SMD can tolerate at most $O(T^{0.95})$ agnostic corruption, which aligns with experiment results. In the second row of Figure D.3, we consider $\alpha=0.1$ for DBGD and $\alpha=0.8$ for NC-SMD. According to the theoretical prediction, both DBGD and NC-SMD can tolerate at most $O(T^{0.9})$ agnostic corruption. From the figure, we can see they have smaller fitted order of regret when $\rho=0.5$ while at a cost of tolerating smaller magnitude of agnostic corruption (it has linear regret when $\rho=0.95$), which reveals intrinsic tradeoff between efficiency and robustness.
Figure D.3: Efficiency-Robustness Tradeoff
D.2 Experiments on Spotify Recommendation Data
Evaluation Setup. We evaluate our approach on Spotify recommendation data (Spotify, 2020). The objective is to recommend songs to incoming users. This dataset includes $17\times 10^{4}$ songs ($|\mathcal{A}|=17\times 10^{4}$), each is described by 15 distinct features (d = 15).
User Profile Identification. In order to model different types of user, we first use Standard Scaler to standardize the dataset. Then we use Kmeans to conduct clustering (Pedregosa et al., 2011).
User Utility. We use the average of the song embedding within each group as the preference vector. Our aim is to recommend songs to users that have the highest cosine similarity (utility function) with their preference vectors. We rescale the cosine similarity to $[-100,100]$.
D.2.1 Performance of Versatile-DB on Spotify Recommendation Data
To test the performance of Versatile-DB, we randomly sample $K=2000$ songs from the Spotify Recommendation Dataset. We identify 2 different user types by clustering the dataset into 2 groups. We use the average of the song embedding within each group as the preference vector and compute the user utility for each recommendation. We consider corruption induced by $\rho$-imperfect user for $\rho\in[0.5,0.75]$. We set $\alpha=0.25$ for DBGD and run both algorithms for $T=100$ iterations. This is because Versatile-DB is extremely computationally intensive and running $K=2000,T=100$ takes around 20 CPU hours. For each user type and for each value of $\rho$, we repeat the experiment over 5 times, each under a different seed. From Figure D.4, we notice that DBGD outperforms Versatile-DB in all $\rho$, which highlights DBGD’s robustness and applicability in real-world data.
Figure D.4: Performance Comparison between DBGD and Versatile-DB

### v2

@@ B Proof for Theorem 2
Since $\lambda\leq\alpha l^{\sigma}_{1}/2$, choosing $\eta_{\rho}=\frac{1}{d}\frac{\sqrt{\log T}}{T^{\rho}}$, we have
$\text{Reg}_{T}$ $\leq 2d\left(C_{0}+\frac{C_{2}}{\sqrt{\log T}}+60C_{\kappa}\sqrt{K}T^{\frac{3/2-\rho}{2^{k+1}}}\right)\sqrt{\log T}T^{\rho}+12Kd\sqrt{\log T}T^{\rho}$
$\leq 144dK\sqrt{\log T}T^{\rho+\frac{3/2-\rho}{2^{k+1}}}.$
$\text{Reg}_{T}\leq\mathcal{O}\left(d\sqrt{\log T}T^{\rho}\right),$
which completes the proof when $\rho\in[0.5,1]$.
C Proof for Proposition 2
Proposition (Proposition 2 restated).
If the total corruption level satisfies $\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|\leq\mathcal{O}(T^{\rho})$, then for any $\alpha\in[\frac{1}{2},1)$, if the utility function is strongly concave, for sufficiently large round $T$, by choosing learning rate $\eta_{\rho}:=\frac{\sqrt{\log T}}{2d}T^{-\alpha}$
  1. 1.
    (Robustness) RoSMID incurs $\text{Reg}_{T}\leq\tilde{\mathcal{O}}(dT^{\alpha}+\sqrt{d}T^{\frac{1}{2}(1-\alpha)+\rho}+dT^{\rho})$.
  2. 2.
    (Efficiency) There exists a strongly concave utility function $\mu$ and a link function $\sigma$ such that $\text{Reg}_{T}$ suffered by RoSMID is at least $\Omega\left(T^{\alpha}\right)$ in scenario without corruption.
C.1 Proof for Robustness Statement in Proposition 2
Proof  From Lemma 4, we know
$\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}$ $\leq 2RdL^{\sigma}\sqrt{\lambda^{*}_{R}+2\phi}T^{\rho}+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq 2RdL^{\sigma}\sqrt{\lambda^{*}_{R}+2\phi}T^{\rho}+2Rd\sqrt{\lambda\eta_{\rho}}L^{\sigma}\sum^{T}_{t=1}\sqrt{t}|c_{t}(a_{t},a^{\prime}_{t})|$
$\leq 2RdL^{\sigma}\sqrt{\lambda^{*}_{R}+2\phi}T^{\rho}+2Rd\sqrt{\lambda\eta_{\rho}T}L^{\sigma}T^{\rho}.$
We get the second inequality by using the fact that the diameter of the action space of $R$. We get the last inequality by using Abel Summation Equation (see Lemma 21). Therefore, by Lemma 3, we have
$\text{Reg}_{T}\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2L^{\mu}L^{\sigma}_{1}R+4RdL^{\sigma}\sqrt{\lambda^{*}_{R}+2\phi}T^{\rho}+4Rd\sqrt{\lambda\eta_{\rho}T}L^{\sigma}T^{\rho}.$
Choosing $\eta_{\rho}=\frac{\sqrt{\log T}}{2d}T^{-\alpha},\alpha\in[0.5,1)$, we have
$\text{Reg}_{T}$ $\leq 2dC_{0}\sqrt{\log T}T^{\alpha}+4d\sqrt{\log T}T^{1-\alpha}+4RdL^{\sigma}\sqrt{\lambda^{*}_{R}+2\phi}T^{\rho}+4R\sqrt{d}L^{\sigma}(\log T)^{0.25}T^{\frac{1}{2}(1-\alpha)}T^{\rho}+2L^{\mu}L^{\sigma}_{1}R$
$\leq O\left(d\sqrt{\log T}T^{\alpha}+\sqrt{d}(\log T)^{\frac{1}{4}}T^{\frac{1}{2}(1-\alpha)}T^{\rho}+dT^{\rho}\right),$
which completes the proof of robustness statement.
C.2 Proof for Efficiency Statement in Proposition 2
Proof  The robustness statement in Proposition 2 implies that RoSMID could afford agnostic corruption level $\mathcal{O}(T^{\frac{1+\alpha}{2}})$. If Lemma 22 holds, this implies that there exists a hard instance which makes RoSMID incur regret in order of $\Omega(T^{\alpha})$ in non-corrupt setting. We can prove this by contradiction. Assume the efficiency statement in Proposition 2 is not true, which implies that for all $\mu$, and $\sigma$, there exists an $\epsilon>0$ such that $\text{Reg}_{T}=c_{0}T^{\alpha-\epsilon}$, for some constant $c_{0}>0$, we assume $\epsilon$ to be the least possible. In this scenario, when we consider a linear link function $\sigma(x)=\frac{1+x}{2}$, Lemma 22 implies that an agnostic corruption budget $2\sqrt{c_{0}}T^{\frac{1+\alpha-\epsilon}{2}}$ suffices to induce linear regret, which contradicts the robustness statement in Proposition 2.
Lemma 22.
For any algorithm $\mathcal{G}$ incurs regret $\text{Reg}_{T}\leq c_{0}T^{\alpha}$, for some constant $c_{0}>0$, on learning from bandit with reward feedback for strongly concave utilities $\mu$ in non-corrupted setting, then there exists an instance $\mu$ such that $\mathcal{G}$ will suffer linear regret under agnostic corruption with budget $\sqrt{c_{0}}T^{\frac{1+\alpha}{2}}$.
Proof  Consider the scenario when $d=2$, the action space $\mathcal{A}:=\{(a_{1},a_{2}):a_{1}\geq 0,a_{2}\geq 0,a_{1}+a_{2}\leq 1\}$, the utility function $\mu_{\theta}(a)=\langle\theta,a\rangle-\frac{1}{2}\|a\|^{2}_{2}$ with $\theta:=[1,0]$. In this scenario, the optimal action is $a^{*}:=[1,0]$. Assume that the adversary want to “fool” the algorithm $\mathcal{G}$ to make it believe the $\theta:=[1,1]$, with optimal arm $a^{*}=[\frac{1}{2},\frac{1}{2}]$, each round, the corruption it should introduce is $c_{t}(a_{t})=a_{t2}$, where $a_{ti}$ is the $i$-th coordinate value of action $a_{t}$. To make reward indistinguishable from these two environments, the total corruption budget which the adversary has to pay is $\sum^{T}_{t=1}|a_{t2}|$. We know that $\mathcal{G}$ incurs regret $\text{Reg}_{T}:=\sum^{T}_{t=1}\langle\theta,a^{*}-a_{t}\rangle-\frac{1}{2}\|a^{*}\|^{2}_{2}+\frac{1}{2}\|a_{t}\|^{2}_{2}=\frac{1}{2}-a_{t1}+\frac{1}{2}a^{2}_{t1}+\frac{1}{2}a^{2}_{t2}\leq c_{0}T^{\alpha}$. Given this constraint, the maximum possible corruption paid by the adversary to make $\theta=[1,1]$ and $\theta=[1,0]$ indistinguishable is $\max\sum^{T}_{t=1}|a_{t2}|=\sqrt{c_{0}}T^{\frac{\alpha+1}{2}}$. Therefore, for any agnostic corruption no less than $\sqrt{c_{0}}T^{\frac{\alpha+1}{2}}$, $\mathcal{G}$ can not distinguish $\theta=[1,0]$ and $\theta=[1,1]$ and has to incur linear regret in either of these environments.
D Proof for Proposition 3:
In this section, we present the regret upper for dueling bandit gradient descent Yue and Joachims (2009) in presence of arbitrary and agnostic corruption. Before presenting the proof, we make several remarks. First, in Yue and Joachims (2009), the utility function $\mu$ is assumed to be strictly concave to ensure the uniqueness of the maximizer. However, in scenarios where the maximizer is unique within the action set for a concave utility function $\mu$, the requirement for “strictness” can be relaxed. Second, we highlight the notation differences. We use $a$ to denote actions while Yue and Joachims (2009) uses $w$. In our proof, we define $P_{t}(a):=\sigma\left(\mu(a_{t})-\mu(a)\right)$ to represent the probability of the event $a_{t}\succ a$. While in Yue and Joachims (2009), it is denoted as $c_{t}(w)$. Just to highlight, its corrupted version $\hat{P}_{t}(a):=\sigma\left(\mu(a_{t})-\mu(a)+c_{t}(a_{t},a)\right)$ is newly introduced in our paper. Moreover, $\bar{P}_{t}(a):=\mathbb{E}_{x\in\mathbb{B}}\left[P_{t}(\mathcal{P}_{\mathcal{A}}(a+\delta x))\right]$ is denoted as $\hat{\epsilon}_{t}(w)+\frac{1}{2}$ in Yue and Joachims (2009). The definition of $\mathcal{F}\left(\mathcal{P}_{\mathcal{A}}\left(a_{t}+\delta u_{t}\right),a_{t}\right)$ is same as $X_{t}\left(\mathcal{P}_{W}(w_{t}+\delta u_{t})\right)$ defined in Yue and Joachims (2009), except different notation. We remark $\mathcal{P}_{\mathcal{A}}(a)$ represent the projection of action $a$ on $\mathcal{A}$.
1 for $t\in[T]$ do
2       Sample unit vector $u_{t}$ uniformly and set $a^{\prime}_{t}=\mathcal{P}_{\mathcal{A}}\left(a_{t}+\delta u_{t}\right)$.
4       Compute gradient $\hat{g}_{t}=-\frac{d}{\delta}\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})u_{t}$.
5       Set learning rate $\eta=\frac{\gamma\delta}{d}$ and update $a_{t+1}=\mathcal{P}_{\mathcal{A}}(a_{t}-\eta\hat{g}_{t})$.
Algorithm 3 Dueling Bandit Gradient Descent (DBGD) (Yue and Joachims, 2009)
Proposition (Proposition 3 restated).
If the total corruption level satisfies $\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|\leq\mathcal{O}(T^{\rho})$, then for any $\alpha\in(0,\frac{1}{4}]$, if utility function $\mu$ is generally concave, for a sufficiently large round $T$, choosing $\gamma:=\frac{R}{\sqrt{T}}$ and $\delta:=\frac{\sqrt{2Rd}}{\sqrt{13L^{\sigma}L^{\mu}}T^{\alpha}}$ for Algorithm 3
  1. 1.
    (Robustness) $\text{Reg}_{T}$ incurred by Algorithm 3 satisfies $\text{Reg}_{T}\leq\mathcal{O}(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha+\rho})$.
  2. 2.
    (Efficiency) There exists a linear utility function $\mu$ and a link function $\sigma$ such that $\text{Reg}_{T}$ suffered by Algorithm 3 is at least $\Omega\left(T^{1-\alpha}\right)$ in scenario without corruption.
D.1 Proof for Robustness Statement in Proposition 3
Proof  The proof of Theorem 3 builds on the proof of Yue and Joachims (2009), where we extend it to a setting in presence of agnostic corruption. The proof follows Step 1-3 introduced in Theorem 2. Similarly, in Step 1, we decompose bias in the gradient caused by corruption. Notice that without corruption, DBGD performs expected gradient ascent, where the gradient $g_{t}$ is defined as follows
In the presence of adversarial corruption, the corrupted gradient is
Therefore, we quantify the bias $\mathbb{E}(b_{t}|a_{t})$ in the following lemma.
Lemma 23.
Proof  (Proof for Lemma 23). Since $a^{\prime}_{t}:=\mathcal{P}_{\mathcal{A}}(a_{t}+\delta u_{t})$, we have
We get the equation (23) by using Lemma 24. We get equation (24) by using Lemma 25.
Lemma 24 (Lemma 2 in Yue and Joachims (2009)).
Fix $\delta>0$, over random unit vector $u$, we have
Lemma 25 (Lemma 1 in Yue and Joachims (2009)).
In Step 2, we decompose regret under dueling bandits into two parts, regret of decision and feedback error.
Lemma 26 (Regret Decomposition for DBGD).
Define $\lambda:=\frac{L^{\sigma}}{L^{\sigma}-\delta L^{\mu}L_{2}}$, $L_{2}$ is the Lipschitz constant for $\sigma^{\prime}$, $\gamma=R/\sqrt{T}$, for $b_{t}$ defined in Lemma 23, we have
$\text{Reg}_{T}$ $\leq\underbrace{\lambda\left(\frac{2Rd\sqrt{T}}{\delta}+13\delta L^{\sigma}L^{\mu}T\right)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Regret\hskip 2.84526ptof\hskip 2.84526ptDecision}}+\underbrace{2\lambda\sum^{T}_{t=1}\mathbb{E}\left(b_{t}^{\top}(a_{t}-a^{*})\right)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Feedback\hskip 2.84526ptError}}.$
Proof  Because of Lemma 27, we have
$\mathbb{E}\left[\sum^{T}_{t=1}\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*})\right]\leq\sum^{T}_{t=1}\mathbb{E}\left(\lambda\nabla\bar{P}_{t}(a_{t})(a_{t}-a^{*})+(3+\lambda)\delta L^{\sigma}L^{\mu}\right)$
$=\lambda\sum^{T}_{t=1}\mathbb{E}(\mathbb{E}(g_{t}|a_{t})^{\top}(a_{t}-a^{*}))+(3+\lambda)\delta L^{\sigma}L^{\mu}T$
$=\lambda\sum^{T}_{t=1}\mathbb{E}(\mathbb{E}(\hat{g}_{t}+b_{t}|a_{t})^{\top}(a_{t}-a^{*}))+(3+\lambda)\delta L^{\sigma}L^{\mu}T$
$=\lambda\sum^{T}_{t=1}\mathbb{E}(\hat{g}_{t}^{\top}(a_{t}-a^{*}))+\lambda\sum^{T}_{t=1}\mathbb{E}(b_{t}^{\top}(a_{t}-a^{*}))+(3+\lambda)\delta L^{\sigma}L^{\mu}T$
Lemma 27 (Lemma 4 in Yue and Joachims (2009)).
Fix $\delta\in(0,\frac{L^{\sigma}}{L^{\mu}L_{2}})$, for $\lambda$ defined in Lemma 26, we have
$\mathbb{E}\left[\sum^{T}_{t=1}\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*})\right]\leq\sum^{T}_{t=1}\mathbb{E}\left(\lambda\nabla\bar{P}_{t}(a_{t})^{\top}(a_{t}-a^{*})+(3+\lambda)\delta L^{\sigma}L^{\mu}\right).$
Noticing that $\|b_{t}\|_{2}\leq\frac{d}{\delta}\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)$. Moreover, $\|a_{t}-a^{*}\|_{2}\leq 2R$. Therefore, in Step 3, we can control the feedback error by the following.
Lemma 28.
$\sum^{T}_{t=1}\mathbb{E}\left(b_{t}^{\top}(a_{t}-a^{*})\right)\leq 2RdL^{\sigma}T^{\rho}/\delta$.
Proof  (Proof for Lemma 28).
$\sum^{T}_{t=1}\mathbb{E}(b_{t}(a_{t}-a^{*}))$ $\leq 2R\frac{d}{\delta}\sum^{T}_{t=1}\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)$
$\leq 2R\frac{dL^{\sigma}}{\delta}\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|$
$=2R\frac{dL^{\sigma}T^{\rho}}{\delta}.$
Lemma 29 (Lemma 3 in Yue and Joachims (2009)).
$\text{Reg}_{T}\leq 2\mathbb{E}\left[\sum^{T}_{t=1}\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*})\right]+5\delta L^{\sigma}L^{\mu}T$
Set $G=\frac{d}{\delta}$, and by Lemma 29, we have
@@ D.1 Proof for Robustness Statement in Proposition 3
$\text{Reg}_{T}\leq\lambda\left(\frac{2Rd\sqrt{T}}{\delta}+13\delta L^{\sigma}L^{\mu}T\right)+4R\frac{\lambda dL^{\sigma}T^{\rho}}{\delta}.$
@@
Therefore, by choosing $\delta:=\frac{\sqrt{2Rd}}{\sqrt{13L^{\sigma}L^{\mu}}T^{\alpha}},\gamma:=\frac{R}{\sqrt{T}}$, and $T>\left(\frac{\sqrt{2Rd}L_{\mu}L_{2}}{\sqrt{13L^{\sigma}L^{\mu}}L^{\sigma}}\right)^{4}$, we have
$\text{Reg}_{T}\leq 2\lambda_{T}\sqrt{26RdL^{\sigma}L^{\mu}}T^{1-\alpha}+2L^{\sigma}\sqrt{26RdL^{\sigma}L^{\mu}}T^{\alpha+\rho}.$
$\lambda_{T}=\frac{L^{\sigma}\sqrt{13L^{\sigma}L^{\mu}}T^{\alpha}}{L^{\sigma}\sqrt{13L^{\sigma}L^{\mu}}T^{\alpha}-L^{\mu}L_{2}\sqrt{2Rd}}$, completes the proof.
D.2 Proof for Efficiency Statement in Proposition 3
Proof  The robustness statement in Proposition 3 implies that DBGD could afford agnostic corruption level $\mathcal{O}(T^{1-\alpha})$. This implies that there exists a hard instance which makes DBGD attain regret in order of $\Omega(T^{1-\alpha})$ in non-corrupt setting, formally described as follows. To start a proof by contradiction, we assume that the efficiency statement in Proposition 3 is false. Specifically, for all problem instance $\mu,\sigma$, there exist a $\epsilon>0$, such that $\text{Reg}_{T}\leq\mathcal{O}(T^{1-\alpha-\epsilon})$. We assume $\epsilon$ to be the least possible. In particular, there exists a pair of instance $\mu,\sigma$ such that $\text{Reg}_{T}=c_{0}T^{1-\alpha-\epsilon}$, $c_{0}$ is a positive constant. In other words, it says that there exists a pair of instance $\mu,\sigma$ such that DBGD suffers regret in $\Theta(T^{1-\alpha-\epsilon})$.
Consider the following problem instance. $\mu(a)=\theta^{\top}a$, specifically $\mu$ is linear function. Let $d=2$ with action set $\mathcal{A}_{1}:=\{(a_{1},a_{2}):a_{1}\geq 0,a_{2}\geq 0,\frac{1}{2}a_{1}+a_{2}-\frac{1}{4}\leq 0\}$ with $\theta=[\frac{1}{2},\frac{1}{2}]$. The optimal arm is $a_{1}=[\frac{1}{2},0]$, which is unique. Let the link function $\sigma(x)=\frac{1}{2}+\frac{1}{2}x$. It is easy to verify that it is rotational symmetric and Lipschitz. Denote $\text{Reg}_{T}$ as the regret occurred by DBGD on this problem instance. We know $\text{Reg}_{T}\leq\mathcal{O}(T^{1-\alpha-\epsilon})$. Consequently, it implies that $a_{2}$ is at most proposed $8c_{0}T^{1-\alpha-\epsilon}$ times. This is because $\text{Reg}_{T}\leq c_{0}T^{1-\alpha-\epsilon}$ and if the proposed action pair $(a_{t},a^{\prime}_{t})$ including $a_{2}$ at round $t$, it occurs regret at least $\frac{1}{8}$.
Now consider a different problem instance with the same $\mu,\sigma$ but different action set, which is $\mathcal{A}_{2}:=\{(a_{1},a_{2}):a_{1}\geq 0,a_{2}\geq 0,\frac{3}{2}a_{1}+a_{2}-\frac{3}{4}\leq 0\}$. The optimal arm in this action set $a_{2}=[0,\frac{3}{4}]$, which is also unique. Proposing arm $a_{1}=[\frac{1}{2},0]$ occurs at least $\frac{1}{8}$ regret. Consider an adversary that pulls the utility of $a_{2}$ from $\frac{3}{8}$ down to $\frac{1}{8}$ whenever it is pulled at a cost of $c_{t}(A_{t})=\frac{1}{4}$. It is equivalent to say that every time when the proposed arm $w$ falls in the region $\mathcal{W}=\{a_{1}\geq 0,a_{2}\geq 0,\frac{1}{2}a_{1}+a_{2}-\frac{1}{4}\geq 0,\frac{3}{2}a_{1}+a_{2}-\frac{3}{4}\leq 0\}$, we assume that the utility is sampled from $\theta^{\top}\mathcal{P}_{\mathcal{A}_{1}}(w)$ and $|c_{t}(w)|=|\theta^{\top}\mathcal{P}_{\mathcal{A}_{1}}(w)-\theta^{\top}w|\leq\frac{1}{4}$. Choosing the total corruption budget as $2c_{0}T^{1-\alpha-\epsilon}$, the adversary can afford to corrupt $8c_{0}T^{1-\alpha-\epsilon}$ times.
This means that as long as the adversary is corrupting, the utility observed in the first problem instance is exactly the same as the utility observed in the second problem instance, in which we pull $a_{2}$ as most $8c_{0}T^{1-\alpha-\epsilon}$. However, $a_{2}$ is the optimal arm in the second problem instance, which implies that the regret occurred on the second problem instance is at least $T-8c_{0}T^{1-\alpha-\epsilon}$, which is linear in $T$. This contradicts the robustness statement in Proposition 3, which says that when agnostic corruption is $2c_{0}T^{1-\alpha-\epsilon}$, the regret upper bound is $O(T^{1-\epsilon})$, which is sublinear. To reconcile the conflicts, we must have the efficiency statement in Proposition 3 true to make the agnostic corruption budget at least $\Omega(T^{1-\alpha})$ to make the parallel world argument valid.
E Additional Experiments
@@
In this section, we list all the experiment details. At the beginning, we introduce the regret order fitting method, and baseline algorithms for comparison.
@@
Baseline Algorithms. We consider three baseline algorithms for comparison, Doubler (Ailon et al., 2014) and Sparring (Ailon et al., 2014)
Doubler is the first approach that transforms a dueling bandit problem into a standard multi-armed bandit (MAB) problem. It operates in epochs of exponentially increasing length: in each epoch, the left arm is sampled from a fixed distribution, and the right arm is selected using a MAB algorithm to minimize regret against the left arm. The feedback received by the MAB algorithm is the number of wins the right arm achieves compared to the left arm. Under linear link assumption, Doubler has been proven to experience regret as the same order as underlying MAB algorithm. For continuous action space and general concave utility, we choose Bandit Gradient Descent (BGD, (Flaxman et al., 2004)), with regret $\mathcal{O}(T^{3/4})$, as the underlying MAB algorithm.
@@
To validate efficiency-robustness trade-off in Proposition 2 and Proposition 3, we try different $\alpha$ values. In the first row of Figure 3, we consider $\alpha=0.05$ for DBGD and $\alpha=0.9$ for RoSMID. We consider $\rho\in[0.5,0.95]$. According to the theoretical prediction, both DBGD and RoSMID can tolerate at most $\mathcal{O}(T^{0.95})$ agnostic corruption, which aligns with experiment results. This is because the estimated order of regret is less than 1. In the second row of Figure 3, we consider $\alpha=0.1$ for DBGD and $\alpha=0.8$ for RoSMID. According to the theoretical prediction, both DBGD and RoSMID can tolerate at most $\mathcal{O}(T^{0.9})$ agnostic corruption. From the figure, we can see they have smaller fitted order of regret when $\rho=0.5$ while at a cost of tolerating smaller magnitude of agnostic corruption (it has linear regret when $\rho=0.95$), which reveals intrinsic tradeoff between efficiency and robustness.
Figure 3: In the first row, we consider $\alpha=0.05$ for DBGD and $\alpha=0.9$ for RoSMID. In the second row, we consider $\alpha=0.1$ for DBGD and $\alpha=0.8$ for RoSMID. Given the $\alpha$, for each algorithm, we tested its performance under $\rho$-Imperfect Human feedback with $\rho=0.5,0.8,0.95$. For each $\rho$, we presented a line plot of the average regret over five simulations, accompanied by $\pm$ one standard deviation shown by the shaded region. In the legend, $o$ denotes the estimated line slope, calculated using least squares on the last $1\%$ of the data.